OpenStudy (anonymous):

how to complete the square of x(x+18)=19

OpenStudy (anonymous):

First distribute, you get x^2 = 18x = 19 To complete the square, you take (1/2) of the coefficient of x, and then square that number. Add that number to both sides of the equation. The coefficient of x is 18, so (1/2) of 18 = 9; and 9^2 = 81. So we add 81 to both sides of the equation. x^2 + 18x + 81 = 19 + 81 The left side is a perfect square.... (x+ 9)^2 = 100 so x+ 9 = +/- sqrt (100) x + 9 = +/- 10 subtract 9 from both sides, you get x = 1 or -19.

OpenStudy (anonymous):

First line should read x^2 + 18x = 19